OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 57x - 8y = 261 -12x + 8y = -36

hero (hero):

Solve each equation for y, set them equal to each other an find x.

OpenStudy (anonymous):

Add the two equations!

OpenStudy (anonymous):

mathematicalpro help

OpenStudy (anonymous):

39x=225 x=5.7

OpenStudy (anonymous):

Solve this for y: -12(5.7)+8y=-36

hero (hero):

Or you could just add the equations.....

OpenStudy (anonymous):

by adding u will get, 45x = 225 x = 225/45 x = 5 -12(5)+8y = -36 -60+8y = -36 8y = -36+60 y = 24/8 y = 3